solve for x in terms of i.
(5+i)x+1-3i=(2-3i)x+2-i

Answer by ewatrrr(24785)


Hi,
solving for x in terms of i
(5+i)x+1-3i=(2-3i)x+2-i
(5+i)x -x(2-3i)= 2-i + 3i - 1
x[(5+i) -(2-3i)] = 1 + 2i
x( 3 + 4i) = (1-2i)
x = (1-2i)/(3+4i)